OpenStudy (anonymous):

Ok now Im stuck lol What is the X=A OR X=B of 4x^2 - 10x - 9 = 0 I got to: (x - 5/4)^2 - 136/16 Would it be X=5/4 OR X= -136/16 Could anyone say if this is right or if I am wrong and give the correct X=A OR X=B? Please?

OpenStudy (phi):

Do you know the quadratic formula? x= (-b +/- sqrt(b*b -4*a*c))/2a

OpenStudy (anonymous):

yh but its a -root which I was told was going the wrong way

OpenStudy (phi):

Oh I see you did complete the square?

OpenStudy (anonymous):

Yh lol That was what the Question asked.

OpenStudy (phi):

I haven't checked your answer, but shouldn't it be an equation: (x - 5/4)^2 - 136/16 = 0

OpenStudy (phi):

If so, move the -136/16 to the right hand side, and then take the square root

OpenStudy (phi):

I checked your work, and where you got -136/16, it should be -61/16 But assuming you fix your mistake, you would have (x-5/4)^2 - 61/16=0 or (x-5/4)^2= 61/16 take the square root of both sides \[x-\frac{5}{4}= \pm\frac{\sqrt{61}}{4}\] and \[x= \frac{5}{4}\pm\frac{\sqrt{61}}{4}\]

OpenStudy (anonymous):

Thanks

OpenStudy (phi):

yw, but I'm still trying to guess how you got 136

OpenStudy (anonymous):

That was the problem D: Ill ask about it tommorow in my lesson.

OpenStudy (anonymous):

I was doing... 4x^2 - 10x - 9 = 0 4(x^2 - 10x/4 - 9/4)=0 therefore if 4(0)=0 x^2 - 10x/4 - 9/4 = 0 (x - 10/4)^2 - 100/16 - 36/16 (x - 10/4)^2 - 136/16 I got the - 100/16 from removing the remainder from ((10/4)x(10/4)) and the 36/16 from 9/4

OpenStudy (anonymous):

oh i see it was (x-5/4)^2??? not (x-10/4) so that would of gave -25/16 - 36/16 = - 61/16 - ahh thats why i got stuck

OpenStudy (anonymous):

Thanks much

OpenStudy (phi):

OK, you have -10x/4 (btw, I would reduce that to -5x/2 ) so you square half of it: -5/4 * -5/4 = -25/16. You squared the whole thing.

OpenStudy (phi):

mystery solved!